# Rollbaccine: Herd Immunity against Storage Rollback Attacks in TEEs (Technical Report)

### v2

@@ Abstract.
copyright: $c$
@@ 1. Introduction
    We introduce Rollbaccine, a device mapper that offers applications rollback resistance (§ 3).
@@
    We provide a formal definition of block device crash consistency (§ 4) and prove that it is preserved by Rollbaccine (Appendix B).
@@
    We show that Rollbaccine adds minimal overhead in most benchmarks and is comparable to state-of-the-art, non-automatic rollback-resistant solutions (§ 7).
@@ 2.1. The Dangers of Rollbacks
Consider for example a TEE application that rate-limits password-guessing attempts (LABEL:lst:password-guessing). It maintains a counter on disk to prevent excessive retries, even across reboots. An attacker could repeatedly crash the TEE and rollback the disk to a state before the counter was incremented, effectively bypassing the guessing limit.
Listing 1: Password-guessing application
with open(counter_file, "r+") as file:
counter = int(file.readline()) + 1
if counter > 10: # Too many attempts
return False
else:
file.write(f"{counter}")
file.flush()
os.fsync(file.fileno())
return pin == real_pin
@@ 3. Towards Rollbaccine
1. Formalisms (§ 4). In order to restore disk to a state it could have recovered to after a benign crash, we need a formal understanding of exactly what states it could have possibly been in. Although file system crash consistency has been extensively studied (ace; crashmonkey; alice; crashMonkeyAndACE; optimisticCrashConsistency; daisyNFS; hydra; chipmunk; yggdrasil; ferrite; monarch; vijayThesis), the field is primarily concerned with avoiding inconsistent states or detecting violations through testing. To the best of our knowledge, no existing work formalizes the correct behaviors for the disk that these file systems rely on.
@@
2. Intercepting (§ 6). Next, we need a well defined interface for intercepting disk operations that is relatively simple to review, maintain, optimize, and trust, as opposed to a custom rollback-resistant application (nimble) or file system. To this end, we implement Rollbaccine as a device-mapper.
@@
Implementing Rollbaccine as a device mapper presents two benefits. First, it is a commonly accepted strategy in industry to augment disk functionality. Dm-crypt (dmCrypt), dm-verity (dmVerity), and dm-integrity (dmIntegrity) are all popular device mappers for enabling disk encryption and (limited) integrity without application modification (awsDmVerity; azureDmCrypt; azureDmVerity; azureStorage; androidDmCrypt; androidDmVerity). Second, device mappers sit below the file system (Figure 1) and are thus file-system agnostic; this allows us to evaluate against both ext4 and xfs in § 7.1 without code modification.
3. Overheads (§ 6). Finally, we must minimize the overhead of replication. Synchronous disk replication on the critical path not only introduces high overhead (§ 7.2), but is also often unnecessary. Existing applications and file systems already carefully engineer their implementation to reduce the number of synchronous writes; these writes are mapped to block I/Os with persistence flags and exposed to the device mapper. It suffices for Rollbaccine to synchronously replicate writes with persistence flags and replicate remaining writes in the background.
@@ 6. Design
Figure 2. Rollbaccine on the critical path with $f=1$.
@@ 6.1.1. Asynchronous writes on the primary
Figure 3. Rollbaccine concurrency handling. “Pop” and “push” are operations over the pending queue, and “insert” and “remove” are operations over the invoked tree.
@@
To address this issue, we impose an ordering on same-block writes by maintaining two data structures: a tree of invoked writes, sorted by write location, and a queue of pending writes, seen in Figure 3. After assigning each write a writeIndexp, the primary atomically checks if it conflicts with any other invoked or pending write. If it does, then the write is placed on the pending write queue and waits to be unblocked. Otherwise, the primary stores its hash, adds the write to the invoked write tree, and submits it to disk. Once the write completes, it is removed from the invoked write tree, and any non-conflicting writes are popped off the pending queue in-order and submitted to disk. At this point, the asynchronous write is marked completed.
@@ 6.1.2. Asynchronous writes arriving at the backups
In order to preserve happens-before relationships between writes to the same block, the backups must still submit writes to disk in order of writeIndexp as assigned by the primary, but do not wait for the disk to finish processing previous writes; only conflicting writes need to block. Concretely, once a backup $b$ receives a write with $\texttt{writeIndex}_{p}=\texttt{writeIndex}_{b}+1$, it atomically increments writeIndexb, then follows the same process depicted in Figure 3. This simultaneously allows non-conflicting writes to be concurrently in-flight while preserving write ordering over individual blocks.
@@ 7. Evaluation
    Generality and Automatability: Can Rollbaccine support unmodified applications, and at what cost? (§ 7.1)
@@
    Performance: How does Rollbaccine compare against non-automatic rollback resistance solutions? (§ 7.2)
@@
    Performance: How do Rollbaccine’s overhead vary as a function of the workload? (§ 7.3)
Implementation. We implemented Rollbaccine as a device mapper for Linux kernel 6.8 available at https://github.com/davidchuyaya/rollbaccine (3,938 LoC). AEAD uses in-kernel AES-GCM; hashing uses HMAC-SHA256. We use in-kernel TCP connections with signed messages for primary-backup communication.
@@ 7.1. Performance Overview
Compared to Unreplicated, DM introduces negligible overhead. Replicated and Rollbaccine respectively reduce throughput by $35\%$ and $15\%$ and increase latency by $54\%$ and $19\%$. This can be attributed to the fact that when benchmarked with TPC-C, roughly every 1 in 5 operations in PostgreSQL are persisted, because every transaction must be durably flushed to PostgreSQL’s Write Ahead Log (WAL) before commit. Both Replicated and Rollbaccine must then synchronously replicate over the network, introducing additional delay, although the latency for Replicated is an order of magnitude greater (§ 7.3). Despite this, the performance penalty is not severe because, at 10 warehouses, TPC-C is contention bottlenecked.
@@ 7.2. Comparison against Nimble
But the culprit is CPU (§ 7.1). Round-trip network latency introduces negligible overhead on the CPU-bottlenecked thread (which is why Rollbaccine and Rollbaccine-sync perform similarly), whereas the asymmetric ECDSA-SHA256 signatures used by Nimble’s messages overwhelm it. NimbleHDFS’s throughput then becomes a function of how many messages it must send; larger batch sizes allow it to amortize signatures across log entries, but once batching is disabled for safety, its throughput reduces to double-digits.
@@ 7.3. Microbenchmarks
Rollbaccine. Reads in Rollbaccine perform similarly to Unreplicated (Figures 6(a), 6(b), 6(c) and 6(d)) because they do not leave the primary, with a maximum of $16\%$ and $21\%$ additional latency and throughput overheads as the result of decryption and maintaining the list of invoked and pending operations; the latter happens in a critical section (§ 6.1). High contention reads (and writes) are the exception (Figure 7), as the conflicting operations are sequentially executed.
@@
In summary, Rollbaccine adds $21\%$ overhead for reads, $45\%$ overhead for asynchronous writes, similar to DM (with the exception of direct writes), and an order of magnitude of overhead for synchronous writes and high contention operations. The under-performance of direct writes and high contention operations is not fundamental; Rollbaccine can be modified to cache and service reads and writes from memory similar to DM. For synchronous writes, Rollbaccine experiences much higher overheads, but, as seen in § 7.1, most applications are designed to use persistence operations sparingly and are minimally affected.
@@ 7.4. Crash Consistency and Recovery
We first break down the performance impact of recovery in Figure 8, plotting time against the number of writes processed by the recovering node. We start with a standard Rollbaccine deployment executing PostgreSQL with TPC-C, as in § 7.1. As it executes, we restart either the primary or backup, overwrite the first 100MB of the 600GB disk to simulate corruption, conduct recovery, then resume TPC-C over the recovered database. Recovery ends after the last shaded region; the following lull in throughput corresponds to TPC-C setup and is present at the beginning of the graph as well. The spikes in throughput are a product of the diverse transactions in TPC-C and are unrelated to recovery.
@@ Appendix B Correctness
Note how the definitions of invocation and responses differ from their definitions in block device crash consistency, which define those operations over the block device (instead of the active primary of Rollbaccine). The active primary in Rollbaccine acts as an additional layer between the application and the block device, delaying invocations to the block device to prevent concurrent accesses to the same blocks (§ 6.1.1), removing read responses with that fail integrity checks (§ 6.1.4), and synchronous write responses until they are replicated (§ 6.1.3).
@@
We know that Rollbaccine processes operations over the same block sequentially based on invocation order (§ 6.1.1), which is unchanged in $\mathcal{S}$. This means that each read must see the previous write, even if the read invocation precedes the write response. This holds despite rollback attacks, because Rollbaccine enforces integrity checks for reads (which would otherwise fail). Since responses immediately follow each invocation in $\mathcal{S}$, each write-read invocation pair satisfies the reads-see-writes precondition and indeed returns the value of the previous write. Therefore $\mathcal{S}$ respects reads-see-writes.

### v3

# Rollbaccine: Herd Immunity against Storage Rollback Attacks in TEEs [Technical Report]

@@ Abstract.
copyright: ccccs: Security and privacy Systems securityccs: Computer systems organization Dependable and fault-tolerant systems and networks
@@ 1. Introduction
    We introduce Rollbaccine, a device mapper that offers applications rollback resistance (Section 3).
@@
    We provide a formal definition of block device crash consistency (Section 4) and prove that it is preserved by Rollbaccine (Appendix B).
@@
    We show that Rollbaccine adds minimal overhead in most benchmarks and is comparable to state-of-the-art, non-automatic rollback-resistant solutions (Section 7).
@@ 2.1. The Dangers of Rollbacks
Consider for example a TEE application that rate-limits password-guessing attempts (Figure 2). It maintains a counter on disk to prevent excessive retries, even across reboots. An attacker could repeatedly crash the TEE and rollback the disk to a state before the counter was incremented, effectively bypassing the guessing limit.
1with open(counter_file, "r+") as file:
2 counter = int(file.readline()) + 1
3 if counter > 10: # Too many attempts
4 return False
5 else:
6 file.write(f"{counter}")
7 file.flush()
8 os.fsync(file.fileno())
9 return pin == real_pin
Figure 2. Password-guessing application
@@ 3. Towards Rollbaccine
1. Formalisms (Section 4). In order to restore disk to a state it could have recovered to after a benign crash, we need a formal understanding of exactly what states it could have possibly been in. Although file system crash consistency has been extensively studied (ace; crashmonkey; alice; crashMonkeyAndACE; optimisticCrashConsistency; daisyNFS; hydra; chipmunk; yggdrasil; ferrite; monarch; vijayThesis), the field is primarily concerned with avoiding inconsistent states or detecting violations through testing. To the best of our knowledge, no existing work formalizes the correct behaviors for the disk that these file systems rely on.
@@
2. Intercepting (Section 6). Next, we need a well defined interface for intercepting disk operations that is relatively simple to review, maintain, optimize, and trust, as opposed to a custom rollback-resistant application (nimble) or file system. To this end, we implement Rollbaccine as a device-mapper.
@@
Implementing Rollbaccine as a device mapper presents two benefits. First, it is a commonly accepted strategy in industry to augment disk functionality. Dm-crypt (dmCrypt), dm-verity (dmVerity), and dm-integrity (dmIntegrity) are all popular device mappers for enabling disk encryption and (limited) integrity without application modification (awsDmVerity; azureDmCrypt; azureDmVerity; azureStorage; androidDmCrypt; androidDmVerity). Second, device mappers sit below the file system (Figure 1) and are thus file-system agnostic; this allows us to evaluate against both ext4 and xfs in Section 7.1 without code modification.
3. Overheads (Section 6). Finally, we must minimize the overhead of replication. Synchronous disk replication on the critical path not only introduces high overhead (Section 7.2), but is also often unnecessary. Existing applications and file systems already carefully engineer their implementation to reduce the number of synchronous writes; these writes are mapped to block I/Os with persistence flags and exposed to the device mapper. It suffices for Rollbaccine to synchronously replicate writes with persistence flags and replicate remaining writes in the background.
@@ 6. Design
((a)) Rollbaccine on the critical path with $f=1$.
((b)) Rollbaccine concurrency handling. “Pop” and “push” are operations over the pending queue, and “insert” and “remove” are operations over the invoked tree.
Figure 3. Rollbaccine’s design.
@@ 6.1.1. Asynchronous writes on the primary
To address this issue, we impose an ordering on same-block writes by maintaining two data structures: a tree of invoked writes, sorted by write location, and a queue of pending writes, seen in Figure 3(b). After assigning each write a writeIndexp, the primary atomically checks if it conflicts with any other invoked or pending write. If it does, then the write is placed on the pending write queue and waits to be unblocked. Otherwise, the primary stores its hash, adds the write to the invoked write tree, and submits it to disk. Once the write completes, it is removed from the invoked write tree, and any non-conflicting writes are popped off the pending queue in-order and submitted to disk. At this point, the asynchronous write is marked completed.
@@ 6.1.2. Asynchronous writes arriving at the backups
In order to preserve happens-before relationships between writes to the same block, the backups must still submit writes to disk in order of writeIndexp as assigned by the primary, but do not wait for the disk to finish processing previous writes; only conflicting writes need to block. Concretely, once a backup $b$ receives a write with $\texttt{writeIndex}_{p}=\texttt{writeIndex}_{b}+1$, it atomically increments writeIndexb, then follows the same process depicted in Figure 3(b). This simultaneously allows non-conflicting writes to be concurrently in-flight while preserving write ordering over individual blocks.
@@ 7. Evaluation
    Generality and Automatability: Can Rollbaccine support unmodified applications, and at what cost? (Section 7.1)
@@
    Performance: How does Rollbaccine compare against non-automatic rollback resistance solutions? (Section 7.2)
@@
    Performance: How do Rollbaccine’s overhead vary as a function of the workload? (Section 7.3)
Implementation. We implemented Rollbaccine as a device mapper for Linux kernel 6.8 available at https://github.com/davidchuyaya/rollbaccine (3,980 LoC). AEAD uses in-kernel AES-GCM; hashing uses HMAC-SHA256. We use in-kernel TCP connections with signed messages for primary-backup communication.
@@ 7.1. Performance Overview
Compared to Unreplicated, DM introduces negligible overhead. Replicated and Rollbaccine respectively reduce throughput by $35\%$ and $15\%$ and increase latency by $54\%$ and $19\%$. This can be attributed to the fact that when benchmarked with TPC-C, roughly every 1 in 5 operations in PostgreSQL are persisted, because every transaction must be durably flushed to PostgreSQL’s Write Ahead Log (WAL) before commit. Both Replicated and Rollbaccine must then synchronously replicate over the network, introducing additional delay, although the latency for Replicated is an order of magnitude greater (Section 7.3). Despite this, the performance penalty is not severe because, at 10 warehouses, TPC-C is contention bottlenecked.
@@ 7.2. Comparison against Nimble
But the culprit is CPU (Section 7.1). Round-trip network latency introduces negligible overhead on the CPU-bottlenecked thread (which is why Rollbaccine and Rollbaccine-sync perform similarly), whereas the asymmetric ECDSA-SHA256 signatures used by Nimble’s messages overwhelm it. NimbleHDFS’s throughput then becomes a function of how many messages it must send; larger batch sizes allow it to amortize signatures across log entries, but once batching is disabled for safety, its throughput reduces to double-digits.
@@ 7.3. Microbenchmarks
Contention is simulated by restricting the set of blocks that each experiment accesses. In the low contention experiments, all threads start execution from the beginning of the block device, leading to an initial spike in contention that tapers off as slightly faster threads no longer contend with slower ones; in the high contention experiments, threads repeatedly access the same initial 4K bytes.
@@
Rollbaccine. Reads in Rollbaccine perform similarly to Unreplicated (Figures 6(a), 6(b), 6(c) and 6(d)) because they do not leave the primary, with a maximum of $16\%$ and $21\%$ additional latency and throughput overheads as the result of decryption and maintaining the list of invoked and pending operations; the latter happens in a critical section (Section 6.1). High contention reads (and writes) are the exception (Figure 7), as the conflicting operations are sequentially executed.
@@
In summary, Rollbaccine adds $21\%$ overhead for reads, $45\%$ overhead for asynchronous writes, similar to DM (with the exception of direct writes), and an order of magnitude of overhead for synchronous writes and high contention operations. The under-performance of direct writes and high contention operations is not fundamental; Rollbaccine can be modified to cache and service reads and writes from memory similar to DM. For synchronous writes, Rollbaccine experiences much higher overheads, but, as seen in Section 7.1, most applications are designed to use persistence operations sparingly and are minimally affected.
@@ 7.4. Crash Consistency and Recovery
We first break down the performance impact of recovery in Figure 8, plotting time against the number of writes processed by the recovering node. We start with a standard Rollbaccine deployment executing PostgreSQL with TPC-C, as in Section 7.1. As it executes, we restart either the primary or backup, overwrite the first 100MB of the 600GB disk to simulate corruption, conduct recovery, then resume TPC-C over the recovered database. Recovery ends after the last shaded region; the following lull in throughput corresponds to TPC-C setup and is present at the beginning of the graph as well. The spikes in throughput are a product of the diverse transactions in TPC-C and are unrelated to recovery.
@@ Appendix B Correctness
Note how the definitions of invocation and responses differ from their definitions in block device crash consistency, which define those operations over the block device (instead of the active primary of Rollbaccine). The active primary in Rollbaccine acts as an additional layer between the application and the block device, delaying invocations to the block device to prevent concurrent accesses to the same blocks (Section 6.1.1), removing read responses with that fail integrity checks (Section 6.1.4), and synchronous write responses until they are replicated (Section 6.1.3).
@@
We know that Rollbaccine processes operations over the same block sequentially based on invocation order (Section 6.1.1), which is unchanged in $\mathcal{S}$. This means that each read must see the previous write, even if the read invocation precedes the write response. This holds despite rollback attacks, because Rollbaccine enforces integrity checks for reads (which would otherwise fail). Since responses immediately follow each invocation in $\mathcal{S}$, each write-read invocation pair satisfies the reads-see-writes precondition and indeed returns the value of the previous write. Therefore $\mathcal{S}$ respects reads-see-writes.